Minutes of the Linac4 Diagnostics Working Group Meeting held on 17 November 2008

Present: U.Raich, J.Tan, G.Bellodi, A.Lokhovitskiy, M.Ludwig, A.Raimondo, C.Dutriat, G.Tranquille, R.Scrivens, A.Rijllart, E.Zorin, T.Hermanns, K.Hanke, B.Mikulec.


1. Communications

2. Follow-up of open actions

3. IPHI diagnostics line - update on status (K.Hanke)

4. Status of the emittance meter application (E.Zorin)

5. New transfer line - status of drawings and open issues (B.Mikulec)

6. AOB

1. Communications

  • The minutes of the last meeting have been approved.

  • K.Hanke presented T.Hermanns, the new fellow in the AB/OP/PSB section, who will work on simulations for the LBE/LBS line upgrade.

  • There will be a meeting on Dec. 1st presenting the LPA (US LHC Accelerator Project) collaboration proposal. The topic on laser profile monitors (LPMs) could be interesting for Linac4, but also for the SPL, as this device would allow to measure the beam profile in a non-intercepting way.

  • K.Hanke reported from the latest 3MeVTCC meeting that the BI racks at the 3 MeV test stand should be filled. U.Raich said that there were 2 BI crates, one of which has already modules installed, but the other one is still empty.

2. Follow-up of open actions

Work for open actions is ongoing.

3. IPHI diagnostics line - update on status

As the IPHI measurement bench will not be delivered, the CERN diagnostics line should already be used for 3 MeV. This is a hot topic as it will already be needed by 2010!

The agreement with CEA is that they would provide some help and hardware. The hardware concerns one spectrometer dipole (without power supply) and 3 (?) quadrupoles (also without power supplies). Magnetic measurements have been done at CEA, which can be consulted in the attached document (Mesures_CEA_ligneIPHI.pdf).
No support can be granted to provide drawings except the drawing for the vacuum chamber (IPL).

Spectrometer design for 3 MeV/12 MeV test bench:
G.Bellodi gave an update on the spectrometer design for 3 MeV/12 MeV (see 3MeVtestbench.pdf). A dipole with magnetic field of 0.17 T (0.34 T) is used as well as two quadrupoles downstream of the dipole and a horizontal SEMgrid at the end of the line before the dump. In front of the dipole there is a slit. For the MEBT commissioning the last quads in the chopper line can be used for matching to a parallel beam onto the slit. However, for the RFQ commissioning an extra doublet will probably be needed before the slit to focalise the beam and to reduce its divergence (max. gradient <20 T/m).

The hardware procurement (at least 1 missing quadrupole and power supplies for dipole plus 4 quadrupoles) has to be followed up!

4. Status of the emittance meter application:

E.Zorin presented the status of the Linac4 emittance meter application (LEM) written in LabView (see LEM.pdf). One already implemented task of the application is the control of the position of the vertical and horizontal SEMgrid and the slit. The application connects to the equipment using the RADE interface library, a FESA class and the equipment driver. The application GUI is also ready, and data from the ADC can be acquired. For the moment only noise data is on-hand. The application is prepared for ppm operation and documentation has been written as well.

Things still to be done: visualisation of motor movement and test with motors, synchronisation test with the DAQ and commissioning of the whole system.

There are several options for data analysis to determine an emittance value. As an EPFL student has already put quite some effort in this analysis, it was proposed to let him continue.

C.Dutriat is working on the cabling for the motor control and potentiometer and expects to be finished by the end of this week. A couple of days should be enough to finalise details on the software side for the motor control. The final hardware equipment is not yet mounted; it was proposed to make the first tests with the slit alone to avoid equipment damage and do the tests with the SEMgrids only once everything else is working.
U.Raich added that 50 ADC cards are currently in production and will be available before the end of the year.

Concern was raised over the leave of E.Zorin in December as this would mean that he will not be available to commission the application. It was suggested to try to find a solution to keep him for a few more months at CERN in order to participate in the emittancs meter commissioning (around March 2009?).
After the meeting an arrangement could be found to keep E.Zorin until 31/03/2009.

5. New transfer line - status of drawings and open issues:

B.Mikulec summarised the outcome of recent meetings and the status on the drawings of the new transfer line between Linac4 and Linac2 (see BI_transferline_update.pdf). A first version of the drawings is already available with only a few open issues. Nevertheless it has to be mentioned that magnet dimensions can change slightly and for most of the BI equipment final dimensions are not yet known. The aim of this presentation was to point out that only for equipment with approx. final dimensions support structures can be designed by a Pakistani team present at CERN from December for 3 months and also later on produced in Pakistan.
U.Raich promised to follow this up with the equipment specialists.

Remark: S.Maury sent out an email after the meeting to inform everybody that the collaboration with the Pakistani team has been cancelled for the moment as we are not yet ready with all aspects of the transfer line. To assure a productive collaboration it is planned to invite the team for January 2010 and plan for a delivery delay of April 2011 for the support structures.

It might be useful to consider adding collimators in the new transfer line after the debuncher to concentrate losses of mainly halo particles (as predicted by simulations). In Linac2 there are carbon rings installed in 2 places for this purpose. Also these 2 rings would have to be exchanged because higher absorption efficiency will be needed with the increased Linac4 beam energy. R.Scrivens mentioned that they have once removed one of the rings; its dimensions corresponded to the available drawings; carbon thickness: 20 mm.

6. AOB

  • Installation of slow wire scanners in chopper line: it would be preferable to install the wire scanners already in advance if available to be able to do the necessary tests and possible modifications. C.Dutriat informed us that RP was pushing for an installation date as late as possible to reduce exposure of the installation team as the wire scanners (which come from the old LEP injector LPI) were still radioactive. This would also mean that the area of the 3 MeV test stand area would have to be classified, which should be avoided in any case. This issue is still open and should be followed up urgently. It was proposed to use the same application as already implemented for ISOLDE (responsibles: J.Sanchez and M.Ludwig).
Assigned to Start date Description State Result
C.Dutriat 2008-11-20 Identify slow wire scanners that can be installed as soon as possible in chopper line. Follow up activation issue with RP and C.Rossi as 3 MeV test stand responsible. edit

  • Faraday Cup readout: The question was asked whether a dedicated application is required for the FC readout or whether an analog signal on a scope would be sufficient. For the initial source tests it was felt that scope acquisition is adequate, but in the long run it would be useful to read out the FC using one ADC channel and a dedicated application. An application is already available in ISOLDE; maybe parts of it could be used. M.Ludwig and A.Lokhovitskiy will look into this.

-- BettinaMikulec - 19 Nov 2008

  • LEM.pdf: Linac4 emittance meter application.

  • BI_transferline_update.pdf: BI equipment in new transfer line.%META:FILEATTACHMENT{name="Mesures_CEA_ligneIPHI.pdf" attachment="Mesures_CEA_ligneIPHI.pdf" attr="" comment="Magnetic measurements of IPHI line magnets." date="1227187523" path="Mesures_CEA_ligneIPHI.pdf" size="55902540" stream="Mesures_CEA_ligneIPHI.pdf" user="Main.BettinaMikulec" version="1"}%
Topic attachments
I Attachment History Action Size Date Who Comment
PDFpdf 3MeVtestbench.pdf r1 manage 107.9 K 2008-11-20 - 14:26 BettinaMikulec Spectrometer for 3 MeV test bench.
PDFpdf BI_transferline_update.pdf r2 r1 manage 450.2 K 2008-11-25 - 14:21 BettinaMikulec BI equipment in new transfer line.
PDFpdf LEM.pdf r1 manage 2339.5 K 2008-11-20 - 15:00 BettinaMikulec Linac4 emittance meter application.
Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r6 - 2010-01-06 - BettinaMikulec
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright